Suppose that a company's annual sales were $1,200,000 in 1999. The annual growth rate of sales from 1999 to 2000 was 16 percent, from 2000 to 2001 it was −5 percent, and from 2001 to 2002 it was 22 percent.

The geometric mean growth rate of sales over this three-year period is calculated as 10.37 percent. Use the geometric mean growth rate and determine the forecasted sales for 2004.
Answer #1
year sales
1999 1200000 0.16
2000 1392000 -0.05
2001 1322400 0.22
2002 1613328

let geometric mean growth rate = r

(1+r)^3 = 1.16 *0.95*1.22

hence

r = 0.1037

sales in 2002 = 1613328

hence forecast in 2004

= 1613328 (1+ r)^2

= 1613328 *1.1037^2

= 1965281